# MMAR: Towards Lossless Multi-Modal Auto-Regressive Probabilistic Modeling

### v2

@@ A.4 CFG With $v$-prediction
A.5 Examples:The effect of the stage 2 training
Figure 5: The impact of the second training stage on image generation capability.
A.6 Examples:Image Generation
@@
Figure 6: Generated images from MMAR-7B
A.7 Detailed Visual Understanding Evaluation Results
@@
A total of 18 visual understanding benchmarks from VLMEvalKit (Duan et al., 2024) are used to evaluate MMAR models comprehensively. The evaluation is also conducted on the existing joint image-text probabilistic models using the publicly available checkpoints [^3] https://huggingface.co/facebook/chameleon-7b [^4] https://huggingface.co/showlab/show-o-w-clip-vit. The detailed evaluation results are shown in table 4. All scores have been scaled to a range of 0 to 100 except that we show the original score of MME benchmarks. The average score is calculated on the normalized score of all the benchmarks including MME.
Table 4: Detailed visual understanding evaluation results.
Benchmark Chameleon-7B Transfusion* Show-o MMAR-0.5B MMAR-0.5B w/ $\epsilon$-pred. MMAR-0.5B w/ VQ MMAR-7B
AI2D 34.81 40.22 32.48 43.43 41.90 41.90 62.63
ChartQA 3.84 9.56 11.32 10.20 10.36 9.36 12.72
DocVQA 1.51 6.72 18.24 7.62 6.77 6.79 10.28
Hallu.Bench 39.01 41.54 40.90 42.80 41.11 41.54 50.47
MathVista 21.90 22.60 23.20 21.60 23.10 22.90 32.10
MMBenchCN 10.14 27.23 0.52 43.99 38.83 31.87 66.49
MMBenchEN 13.32 29.47 42.44 48.45 45.53 37.54 66.32
MMEP 125.8 594.3 1182.7 882.1 880.7 618.2 1393.9
MMEC 33.9 206.1 225.0 256.8 232.1 273.2 320.0
MMMU 24.00 29.33 26.44 29.33 25.33 29.67 42.78
MMStar 20.47 28.13 32.00 32.13 31.07 28.07 40.00
MMVet 7.34 13.90 20.87 18.49 17.98 14.45 27.80
OCRBench 0.50 2.30 15.20 18.70 7.10 2.10 23.10
POPE 30.86 66.90 84.50 70.74 71.14 66.98 83.02
RealWorldQA 27.06 36.99 27.97 38.30 35.16 36.60 52.42
ScienceQA 44.83 45.92 41.82 47.54 45.21 45.35 71.05
SEEDBench 34.61 42.40 51.61 55.70 53.72 44.93 64.52
TextVQA 5.43 9.94 38.35 16.77 12.40 9.46 22.05
Average 18.34 28.26 33.06 34.56 32.21 29.70 46.52
A.8 Analysis on the Visual Understanding Ability of Transfusion-Like Methods
In this section, we will quantitatively analyze the incomplete information usage problem in transfusion-like methods.
The experiments are conducted using our re-implemented transfusion model. The model is evaluated on MMBench using different noise-level image input. Specifically, we take the diffusion time-step of 0, 125, 250, 375, 500, and 750 for testing. Among the MMBench tasks, we focus on tasks whose performance is significantly affected by noise-levels. Fig. 7 depict the performance trend of these tasks across different noise levels.
Figure 7: The trend of MMBench tasks’ performance with respect to input image noise levels. (A): The best performance of most tasks are achieved at a non-zero noise level. (B): The object localization performance consistently gets higher as the noise level grows up.
Intuitively, increasing noise levels should lead to a consistent performance decline across all visual understanding tasks due to information loss. However, our findings reveal a more nuanced relationship. We find that the best performance of most visual understanding tasks are achieved at a non-zero noise level. Especially, the object localization task, which is more focused on the image layout information, consistently performs better as the noise level grows higher. This suggests that transfusion-like methods, when limited to a single noise level input, fail to fully exploit the information of the input image. Instead, visual information appears to be distributed across different noise levels within the model, with the input noise level dictating which information is ultimately utilized.

### v3

@@ 6.4 CFG With $v$-prediction
Figure 9: Generated images from MMAR-7B after the third training stage (1/2).
6.5 Impact of CFG Scaling
We select models from the second and fourth epochs of the first stage as starting points for the second stage, train them for 3 epochs, and then test the MSCOCO FID-30K under varying CFG intensities. As shown in Fig. 8, our method achieves better FID scores as the CFG scale increases from 1 to 10. It is worth noting that most probabilistic generative models typically have a CFG scale between 1.5 and 5. Additionally, it is observed that a longer training duration in the first stage consistently results in better generation outcomes at all CFG scales.
6.6 Detailed Visual Understanding Evaluation Results
@@
A total of 18 visual understanding benchmarks from VLMEvalKit [12] are used to evaluate MMAR models comprehensively. The evaluation is also conducted on the existing joint image-text probabilistic models using the publicly available checkpoints [^3] https://huggingface.co/facebook/chameleon-7b [^4] https://huggingface.co/showlab/show-o-w-clip-vit. The detailed evaluation results are shown in Table 6. All scores have been scaled to a range of 0 to 100 except that we show the original score of MME benchmarks. The average score is calculated on the normalized score of all the benchmarks including MME.
6.7 Examples:Image Generation Sample
In Fig.910, we showcase some generated examples from MMAR-7B after the third stage training, featuring animals, plants, real-world scenes, artistic scenes, and counterfactual themes.
Figure 10: Generated images from MMAR-7B after the third training stage (2/2).
6.8 Numerical Error Empirical Validation
Figure 11: Diffusion sampling results with different prediction targets and training precision. Figures are scaled differently to include all generated samples.
We provide direct empirical demonstration using a 2D chessboard distribution (class-conditioned quadrant mapping). As Fig.11 reveals: (1) v-prediction maintains stability even with bf16 precision at scale (1M samples, Fig.(a)), achieving $<10^{-6}$ token error rate. (2) $\epsilon$-prediction exhibits visible artifacts (Fig.(b),(c)) with 0.2% token error rate (bf16), equivalent to 51.2% image-level failure for 256-token images. (3) Precision elevation (fp32) reduces artifacts to 0.04%. These observations validate v-prediction’s critical role in mitigating low-precision training risks.